# Explainable Anomaly Detection in Images and Videos: A Survey

### v2

@@ V-B Video Anomaly Detection
Street Scene [122] 2020 Se 1 203,257 56,847 146,410 17 Jaywalking, Biker outside lane, Loitering, Dog on sidewalk, Car outside lane, Worker in bushes, Biker on sidewalk, Pedestrian reverses direction, Car u-turn, Car illegally parked, Person opening trunk, Person exits car on street, Skateboarder in bike lane, Person sitting on bench, Metermaid ticketing car, Car turning from parking space, Motorcycle drives onto sidewalk
TAD [123] 2021 We - 540,272 452,220 88,052 7 Vehicle accidents, Illegal turns, Illegal occupations, Retrograde motion, Pedestrian on road, Road spills, The else
UBnormal [124] 2022 Su 29 236,902 116,087 92,640 22 Running, Falling, Fighting, Sleeping, Crawling, Having a seizure, Laying down, Dancing, Stealing, Rotating $360$ degrees, Shuffling, Walking injured, Walking drunk, Stumbling walk, People and car accident, Car crash, Running injured, Fire, Smoke, Jaywalking, Driving outside lane, Jumping
NWPU Campus [125] 2023 Se 43 1,466,073 1,082,014 384,059 28 Climbing fence, Car crossing square, Cycling on footpath, Kicking trash can, Jaywalking, Snatching bag, Crossing lawn, Wrong turn, Cycling on square, Chasing, Loitering, Scuffle, Littering, Forgetting backpack, U-turn, Battering, Driving on wrong side, Falling, Suddenly stopping cycling in the middle of the road, Group conflict, Climbing tree, Stealing, Illegal parking, Trucks, Protest, Playing with water, Photographing in restricted area, Dogs
In this section, we introduce some popular datasets for video anomaly detection and give some insight on whether it can be applied to different explainable methods. The detailed statistics of the listed datasets are given in Tab. V, which, for the first time, lists the anomaly classes of each dataset for a better understanding of the dataset’s nature.
UMN [116]
contains a series of video clips that are recorded from different scenes, such as an indoor hallway, an outdoor lawn, and a plaza. The abnormal events are characterized by people suddenly starting to run in all directions, simulating an escape or panic scenario. However, it does not provide object-level anomaly labels, which limits its utility in evaluating explainable video anomaly detection models.
UCSD Pedestrian 1 (Ped1) [117]
includes video segments showing crowds moving and walking. It comprises 34 training video samples and 36 testing video samples. Additionally, a subset of 10 Ped1 clips is provided with hand-crafted binary masks at the pixel level, which serve to pinpoint areas that contain anomalies. This feature enables the evaluation of an algorithm’s ability to localize anomalies. The inclusion of mask regions facilitates the assessment of explainability through the anomaly localization task.
UCSD Pedestrian 2 (Ped2) [117]
contains $16$ training and $12$ test videos. It is meticulously crafted to comprehensively encompass a broad spectrum of anomalous activities such as unauthorized vehicles, bicycles, skateboarders, and wheelchairs intruding into pedestrian regions. Notably, these datasets are celebrated for not only their complexity but also for embodying a wide range of real-world scenarios, thereby serving as a robust platform for the explainability of advanced anomaly detection algorithms, and catalyzing research towards innovative solutions for recognizing and addressing anomalies in dynamic, pedestrian-rich environments. The dataset is very suitable for attention-based explainable method, but due to the lack of detailed attributes, it cannot be applied to the attribute-based explainable method.
Subway [115]
encompasses two videos, including a $96$-minute entrance gate segment and a $43$-minute exit gate clip. It offers a rich tableau of both regular and irregular passenger behaviors within a bustling subway environment. Anomalies are typified by passengers moving in erroneous directions and ticket evasion attempts, presenting a robust and challenging testing ground for the creation and assessment of anomaly detection algorithms in real-world, high-density transit scenarios. The dataset is very suitable for attention-based explainable methods but similar to Ped2, it cannot give enough information to evaluate attribute-based explainable methods.
Avenue (Ave) [118]
offers a discerning collection of videos specifically curated for anomaly detection research. Constituting videos primarily focused on avenues, the dataset provides researchers with a variety of challenging scenarios under diverse environmental and lighting conditions. It uniquely encompasses a wide range of anomalous events, such as loitering, throwing objects, and running, amidst everyday pedestrian activities. Captured via fixed-position cameras, the dataset facilitates explainability evaluations by providing a balanced mixture of ordinary and anomalous events. Similar to Ped2, it can only work for the Attention-based explainable method.
Street Scene (SST) [122]
has $46$ training and $35$ testing video sequences, all captured at a high resolution of $1280\times 720$, observing a bustling two-lane street scene during daytime hours. With an enriched tapestry of activities, including cars navigating, and pedestrians in varied motions, the dataset introduces multifaceted challenges for anomaly detection. Remarkably, Street Scene contains $205$ naturally occurring anomalous events, which span illegal activities such as jaywalking and unauthorized U-turns, infrequent occurrences like pet-walking and meter maid interventions, thereby serving as a comprehensive, real-world benchmark for developing and evaluating anomaly detection models amidst the complexity and variability of urban environments.
UCF-Crime (UCF) [120]
represents a substantial contribution to the domain of anomaly detection, boasting $1,900$ videos that encompass 13 distinct anomalous event categories, recorded via indoor and outdoor CCTV cameras under varied lighting conditions. Categories encompass a wide spectrum of unlawful and disruptive activities including, but not limited to, Abuse, Arrest, Assaults, and Burglary, thereby providing a comprehensive and challenging framework for the development and empirical evaluation of video anomaly detection algorithms amidst a variety of complex, real-world scenarios. This provides a rich context for models to show their explainability. This dataset provides detailed anomaly labels which is useful in almost all of the existing explainable methods, including reasoning-based methods, and attribute-based methods.
ShanghaiTech Campus (STC) [119]
includes a varied compilation of $437$ videos, with durations spanning $15$ seconds to over a minute, recorded under an array of circumstances and illumination conditions in outdoor settings via CCTV cameras. Notably, it presents multifaceted challenges through its inclusion of complex lighting situations and multiple camera angles, thereby establishing a comprehensive platform for researchers to delve into, evaluate, and develop anomaly detection algorithms within contextually rich and diversely illuminated environments. This dataset provides detailed actions and complex anomalies, so some explainable techniques like scene graph can work well in this dataset, which is suitable for attention-based method and reasoning-based methods.
NWPU Campus [125]
is proposed for semi-supervised video anomaly detection. It stands as the largest and most complex dataset in its field, encompassing $43$ scenes, $28$ classes of anomalous events, and a total of $16$ hours of video footage. Notably, it includes scene-dependent anomalies, indicating that an event might be considered normal in one scene but abnormal in another. It is suitable for interpreting VAD models in complex scenarios. Consequently, this characteristic presents a challenging problem for developing scene-specific explainable methods.
XD-violence (XD) [121]
focuses on recognizing and understanding violent activities. This dataset is meticulously compiled, comprising a diverse array of videos that encapsulate various facets and degrees of violent actions in numerous settings and contexts. Encompassing scenarios that range from mild altercations to more severe and explicit forms of violence, the XD-violence dataset poses a complex challenge for video anomaly detection models. It not only serves as a comprehensive benchmark for evaluating the efficacy of existing models but also catalyzes the advancement of novel methods tailored to discern and categorize different intensities and natures of violent activities amidst multifarious environmental variables. This dataset has comprehensive anomaly events and labels which make it suitable for most of the explaianble methods.
X-MAN [68]
is derived and reannotated from notable datasets like ShanghaiTech, Avenue, and UCSD Ped2. Each frame is meticulously annotated with between $1$ and $5$ explanation labels, each elucidating a distinct rationale for the frame’s anomaly, thereby catering to a multitude of anomalous events within a singular frame—such as simultaneous occurrences of running and bike-riding. Cumulatively, the dataset gives $40,618$ labels across all frames, with a predominant $22,640$ anomalies attributed to actions, followed by $14,828$ ascribed to anomalous objects, while the residual pertains to anomalies related to location. Significantly, the dataset delineates $42$ anomalous actions and $13$ anomalous objects, presenting a comprehensive, multi-faceted benchmark for developing and assessing anomaly detection algorithms in varied contexts. It is designed for evaluating explainable methods as it provides very detailed anomaly reasons, so it is very suitable for attributes-based methods.
UBnormal [124]
is an open-set, supervised collection of various virtual scenes designed for video anomaly detection. It stands out from current datasets by providing pixel-level annotations of abnormal events during training, which allows for the employment of fully-supervised learning techniques in detecting unusual events for the first time. To preserve the open-set structure, the developers have carefully arranged for different types of anomalies to be present in the training and testing video sets. This dataset provides pixel-level anomaly labels, so it can be applied to evaluate explainability, such as localization, for the explainable VAD method.
TAD [123]
focuses on traffic anomalies, differing significantly from other anomaly detection benchmarks by covering real-world traffic conditions. It aims to enhance the diversity and complexity of anomaly detection scenarios, particularly in traffic monitoring for early warning and emergency assistance, by including various real-world anomalies across different conditions and times of day. Since existing anomaly detection methods primarily focus on human behavior, the traffic anomaly dataset bridges the gap for traffic anomaly detection scenarios. Furthermore, it provides object-level labels and annotations for anomalous events, which makes quantifying explainability possible.
@@ VI Evaluation Metrics
VI-A For images
Image-level metrics
For image-level anomaly detection, i.e., determining whether the sample is anomalous or not, the most prominent metrics are the Area Under the Receiver Operating Characteristic (AUROC) and the Area Under the Precision-Recall curve (AUPR). The ROC curve is a visual plot that demonstrates the performance of a binary classification system across a range of decision thresholds. It maps the rate of true positives (TPR) to the rate of false positives (FPR) on a graph. The AUROC quantifies the overall ability of the classifier to distinguish between the positive and negative classes. An anomaly detector with An AUROC score of 0.5 suggests that the model is random guessing, whereas a score of $1$ signifies an impeccable anomaly detector. Precision-recall curves are particularly beneficial for assessing classification in imbalanced datasets. The AUPR offers a summarized measure of the classifier’s performance across different thresholds. For datasets with a significantly low anomaly ratio, the AUPR can be more informative than the AUROC.
Pixel-level metrics
For pixel-level anomaly localization and segmentation of images, the most commonly adopted evaluation metrics are AUROC and IoU (Intersection over Union). Assessing the effectiveness of anomaly segmentation techniques at the pixel-by-pixel scale considers the result of classifying each pixel to be of equal significance. Just as image-level anomaly detection, a pixel’s classification can also fall into one of four categories: TP, TN, FP, or FN. For each of these scenarios, the aggregate count of pixels from the test dataset is determined as:
TP $=\sum_{i=1}^{n}\left|\{p\mid L_{i}(p)=1\}\cap\{p\mid S_{i}(p)>t\}\right|,$ (9)
TN $=\sum_{i=1}^{n}\left|\{p\mid L_{i}(p)=0\}\cap\{p\mid S_{i}(p)\leq t\}\right|,$ (10)
FP $=\sum_{i=1}^{n}\left|\{p\mid L_{i}(p)=0\}\cap\{p\mid S_{i}(p)>t\}\right|,$ (11)
FN $=\sum_{i=1}^{n}\left|\{p\mid L_{i}(p)=1\}\cap\{p\mid S_{i}(p)\leq t\}\right|,$ (12)
[…]
While several metrics exist for evaluating segmentation algorithms, the Intersection over Union (IoU) metric stands out for its efficacy. Specifically, within the realm of anomaly segmentation, the set of anomalous predictions is denoted as $P=\bigcup_{i=1}^{n}\{p|S_{i}(p)>t\}$, and the set of ground truth anomalous pixels is represented as $G=\bigcup_{i=1}^{n}\{p|L_{i}(p)=1\}$. In parallel to the aforementioned relative metrics, the IoU for the ’anomalous’ category can be articulated using absolute pixel classification metrics as:
$\text{IoU}=\frac{|P\cap G|}{|P\cup G|}=\frac{\text{TP}}{\text{TP}+\text{FP}+\text{FN}}.$ (13)
While these metrics are computationally straightforward, their pixel-independent nature can bias results toward larger anomalous regions [25, 96]. Detecting one extensive defect might offset multiple undetected minor defects. Given the variance in defect sizes across the proposed dataset, it’s crucial to consider metrics based on each ground truth’s connected component, which leads to Region-level metrics.
Region-level metrics
Rather than assessing each pixel individually, metrics at the region level take an average performance across each contiguous segment of the actual data. This method is particularly beneficial when recognizing tiny irregularities is deemed as crucial as spotting larger ones, a scenario commonly found in real-world applications. The most commonly used region-level metric in anomaly segmentation is the per-region overlap (PRO) metric. To start, each test image’s actual data is broken down into its connected segments. Suppose $A_{i,j}$ represents the group of pixels identified as anomalies for a particular segment $j$ in the authentic image $i$, and $G_{i}$ stands for the group of pixels anticipated to be anomalous at a set threshold t. The PRO can subsequently be calculated as
$\text{PRO}=\frac{1}{N}\sum_{i}\sum_{k}\frac{|G_{i}\cap A_{i,j}|}{A_{i,j}},$ (14)
where $N$ represents the total count of true elements in the dataset under evaluation. The PRO metric bears a strong resemblance to the TPR. The key distinction is that, with PRO, the TPR is averaged across each ground truth region rather than across every individual pixel in the image.
VI-B For videos:
Besides the AUC that is used in both image video anomaly detection and video anomaly detection, there are some unique metrics for videos.
Track-Based Detection Rate [122]
In the realm of video anomaly detection, where accurate identification and classification of anomalous activities are paramount, the Track-Based Detection Rate (TBDR) metric emerges as a pivotal criterion, meticulously assessing detection performance across video sequences. The TBDR quantitatively juxtaposes the number of accurately detected anomalous tracks against the total number of anomalous tracks present, as expressed in
$TBDR=\frac{\text{num. of anomalous tracks detected}}{\text{total num. of anomalous tracks}}$ (15)
An anomalous track is deemed detected if at least a fraction $\alpha$ of the ground truth regions within the track is identified, and a ground truth region in a frame is considered detected if the Intersection Over Union (IOU) with a detected region is at least $\beta$. Concurrently, the False Positive Rate per frame (FPR), expressed as
$FPR=\frac{\text{total false positive regions}}{\text{total frames}}$ (16)
It underscores the model’s precision by quantifying the number of falsely identified regions relative to total frames, whereby a detected region is labeled as a false positive if the IOU with every ground truth region in the frame is $<\beta$. Significantly, while a singular detected region might encompass two or more ground truth regions, ensuring each truth region is detected, such occurrences are sporadic, thereby entrenching TBDR as a vital metric in distilling the efficacy and precision of anomaly detection models within complex video datasets.
RBDR: Region-Based Detection Criterion [122]
The RBDR is determined across all frames in a testset in relation to the quantity of false positive regions for each frame, mathematically expressed as
$RBDR=\frac{\text{num. of anomalous regions detected}}{\text{total num. of anomalous regions}}$ (17)
It assesses detection capabilities by appraising the number of precisely detected anomalous regions about the entirety of anomalous regions present. Similar to the track-based criterion, a region of ground truth is considered successfully detected if the overlap measure, known as the Intersection Over Union (IOU), between the actual ground truth and the detected region is at least $\beta$. The RBDR considers all ground truth anomalous regions throughout the entirety of frames within a test set. Simultaneously, the number of false positives per frame is evaluated similarly to the track-based detection criterion. Embarking on detection criteria inevitably lands on the balance between detection rate (true positive rate) and false positive rate, a relationship which can be elucidated through a Receiver Operating Characteristic (ROC) curve, deduced by varying the threshold on the anomaly score dictating the detection of anomalous regions. When a singular evaluative figure is sought, the performance is adeptly summarized utilizing the average detection rate for false positive rates ranging from 0 to 1, i.e., the area under the ROC curve for false positive rates $\leq 1$.
VII Future Directions
VII-A Quantifying explainable methods.
In the VAD task, there are plenty of explainable methods that give explanations from every aspect, including pixel level, object level, and so on. However, none of them has quantified their explanation performance. [68] tried to quantify the explanation performance by showing the anomaly classification performance, but other methods can barely follow due to the different explanation directions. In this case, rather than showing some selected examples to prove the effectiveness of one method, it is better to quantify the effectiveness of an explainable method. For example, if a method explains by providing an attention map to illustrate the explanation, it is preferred to use the Intersection Over Union metric like the image segmentation task. A dataset with such ground truths is also needed to achieve the target. In addtion, existing metrics in explainable AI can be adapted for anomaly explanation, including those using feature importance (e.g., Saliency [126], Gradient SHAP [127], Deepflit [128], Representation latent shift [129], etc.) and example importance (e.g., Influence function [130], Tracing gradient descent [131], Importance score [129], etc. ).
VII-B Explaining the anomaly detection methods involving self-supervised learning.
Current state-of-the-art 2D anomaly detection methods, either for the image or video data, are more or less related to self-supervised learning [21, 5, 20]. However, there exist few works clearly explaining why creating pretext tasks can improve anomaly detection performance. The main difficulty in interpreting the self-supervised anomaly detection techniques lies in that it is hard to explain the performance difference between the normal and abnormal samples on the auxiliary task. To investigate the explainability of such methods, researchers may need to refer to the theory of self-supervision and data augmentation.
VII-C Building new benchmarks for explainable 2D anomaly detection.
Existing image and video anomaly detection benchmarks [25, 26, 27, 8] are still too simple and ideal for real-world applications. One unavoidable and important issue is that normal objects or behaviors in one background might be normal in another. For instance, two people punching each other is considered normal in the boxing ring, while abnormal in the classroom. Therefore new benchmarks are needed to handle such deceptive cases and the desired methods are supposed to be able to conduct reasoning among the background and objects with sensible explanations.
VII-D Foundation models for 2D anomaly detection
Foundation models or large pre-trained models, such as GPT-4 [132], LLaVA [133], SAM [134], and DINO [135], have demonstrated exceptional capabilities in reasoning and understanding images. These models are particularly powerful due to their pre-training on vast datasets comprising images and their associated textual descriptions. Consequently, they possess the ability to discern normal and abnormal objects when presented with either images or text. This intrinsic knowledge proves invaluable for various downstream tasks, including image anomaly detection and video anomaly detection. For instance, AnomalyGPT [54] leveraged the large multi-modal model, pandaGPT [136], to detect anomalies in images. By simply inputting an image to pandaGPT and asking whether anomalies are present, their method not only identifies anomalies but also provides explanations, such as identifying cracks or cuts. On the other hand, SAA [53] localizes anomalies without the need for explicit training. They achieve this by adapting existing foundation models, DINO [135] and SAM [134], through hybrid prompt regularization. These successful applications of foundation models in image anomaly detection underscore their robust anomaly detection capabilities. However, current methods harness only a fraction of the potential offered by foundation models, such as their zero-shot and few-shot abilities. The extensive reasoning capabilities of these models remain largely unexplored. Notably, there has been a lack of research applying foundation models to video anomaly detection, primarily due to the complexity of video anomalies. In video anomaly detection, anomalies encompass multiple dimensions, including object anomalies, motion anomalies, and even background anomalies. Simple retrieval methods are often insufficient for this task. However, with the aid of their advanced reasoning abilities, foundation models can easily discern whether a particular motion is plausible within a given background context and provide explanations. Moreover, foundation models offer a promising approach for adapting to various scenes. For instance, while street fighting should be classified as an anomaly, it is normal within an octagonal cage context. Traditional video anomaly detection models struggle to distinguish between these scenarios, often leading to incorrect predictions. In contrast, foundation models can effectively handle such distinctions due to their ability to capture scene information through their extensive parameterization, enabling them to treat different scenes separately.

### v4

@@ V-B Video Anomaly Detection Datasets
Street Scene [169] 2020 Se 1 203,257 56,847 146,410 17
Jaywalking, Biker outside lane,
Loitering, Dog on sidewalk, Car outside lane,
Worker in bushes, Biker on sidewalk, Pedestrian reverses direction,
Car u-turn, Car illegally parked, Person opening trunk,
Person exits car on street, Skateboarder in bike lane,
Person sitting on bench, Metermaid ticketing car,
Car turning from parking space, Motorcycle drives onto sidewalk
TAD [170] 2021 We - 540,272 452,220 88,052 7
Vehicle accidents, Illegal turns, Illegal occupations,
Retrograde motion, Pedestrian on road, Road spills, The else
UBnormal [171] 2022 Su 29 236,902 116,087 92,640 22
Running, Falling, Fighting, Sleeping,
Crawling, Having a seizure, Laying down, Dancing,
Stealing, Rotating 360360 degrees, Shuffling, Walking injured,
Walking drunk, Stumbling walk, People and car accident,
Car crash, Running injured, Fire, Smoke,
Jaywalking, Driving outside lane, Jumping
NWPU Campus [172] 2023 Se 43 1,466,073 1,082,014 384,059 28
Climbing fence, Car crossing square, Cycling on footpath,
Kicking trash can, Jaywalking, Snatching bag, Crossing lawn,
Wrong turn, Cycling on square, Chasing, Loitering,
Scuffle, Littering, Forgetting backpack, U-turn,
Battering, Driving on wrong side, Falling,
Suddenly stopping cycling in the middle of the road, Group conflict,
Climbing tree, Stealing, Illegal parking, Trucks, Protest,
Playing with water, Photographing in restricted area, Dogs
The datasets commonly used for VAD are summarized in Tab. IV, which includes details on settings, frame counts, anomaly classes, and relevance to explainability evaluation. Below, we briefly describe some key datasets and their implications for explainable VAD.
Early Benchmarks (UMN [163], Ped1/Ped2 [164], Subway [162], Avenue [165])
These datasets feature relatively simple scenes and anomalies (e.g., running crowds, vehicles in pedestrian zones, wrong direction). They are useful for evaluating basic spatio-temporal localization explanations (e.g., attention maps, reconstruction errors). Ped1 provides pixel masks for some clips, allowing direct evaluation of localization accuracy. However, the limited complexity might not challenge more sophisticated reasoning or attribute-based explanation methods.
Complex Scene Benchmarks (ShanghaiTech [166], Street Scene [169], NWPU Campus [172])
These datasets offer more diverse scenes, camera angles, lighting conditions, and a wider range of anomaly types, including complex interactions and subtle events. They are better suited for evaluating advanced explanation methods:
  • STC Its diversity supports reasoning-based explanations (e.g., scene graphs capturing interactions). The X-MAN re-annotation [97] provides explicit reason labels, directly facilitating evaluation of attribute-based explanations.
  • Street Scene Features numerous, varied, naturally occurring anomalies. Good for evaluating fine-grained localization and explanations related to specific actions or interactions.
  • NWPU Campus Very large-scale with many scenes and anomaly types, including scene-dependent anomalies (e.g., fighting is normal in a boxing ring, abnormal elsewhere). This explicitly challenges methods to provide context-aware explanations, likely requiring reasoning capabilities.
Weakly-Supervised Benchmarks (UCF-Crime [167], XD-Violence [168], TAD [170])
These datasets typically provide only video-level labels (indicating if an anomaly occurs in a long video, often without precise timing or location). Explaining anomalies here is challenging, often requiring methods that can link coarse labels to specific spatio-temporal segments (e.g., Multiple Instance Learning approaches). Generating fine-grained explanations like pixel maps or detailed attributes is difficult due to the lack of precise ground truth.
  • UCF-Crime/XD-Violence Focus on real-world criminal/violent events.
  • TAD Focuses on traffic anomalies, providing object-level labels for evaluating explanations in this specific domain.
Specialized/Supervised Benchmarks (X-MAN [97], UBnormal [171])
These datasets are specifically designed with explainability or detailed supervision in mind.
  • X-MAN As mentioned, provides multiple textual reason labels per anomalous frame (action, object, location), directly targeting the evaluation of attribute-based explanation methods.
  • UBnormal Provides pixel-level annotations in a supervised setting (anomalies seen during training, but test set is open-set). Useful for evaluating precise localization explanations and supervised attribute learning in a controlled (virtual) environment.
Overall, the choice of dataset significantly influences the type of explainable method that can be developed and evaluated. Datasets with fine-grained annotations (pixel masks, bounding boxes, attribute labels) are essential for quantitatively evaluating localization and attribute-based explanations. More complex datasets with diverse scenes and context-dependent anomalies are needed to push research towards more robust reasoning-based explanations.
@@ VI Evaluation Metrics
Evaluating explainable anomaly detection involves assessing both the accuracy of the anomaly detection itself and the quality of the generated explanations. Standard metrics focus on detection/localization performance, while evaluating explanations requires considering aspects like faithfulness, stability, and utility.
VI-A Metrics for Anomaly Detection Performance
These metrics quantify how well the system identifies and localizes anomalies, irrespective of the explanation provided.
Image-level / Frame-level Detection
For tasks determining whether an entire image or video frame is anomalous, standard binary classification metrics are used:
  • Area Under the Receiver Operating Characteristic (AUROC) Plots True Positive Rate (TPR) vs. False Positive Rate (FPR) across different thresholds. A score of 1 indicates perfect separation, 0.5 indicates random guessing.
  • Area Under the Precision-Recall curve (AUPR) Plots Precision vs. Recall. More informative than AUROC for highly imbalanced datasets typical in anomaly detection.
  • F1-Score The harmonic mean of Precision and Recall at a specific threshold.
Pixel-level / Region-level Localization
For tasks requiring localization of anomalous regions within images or video frames:
  • Pixel-wise AUROC/AUPR Calculated by treating each pixel as an individual classification instance. Can be biased towards larger anomalous regions [39].
  • Intersection over Union (IoU) Measures the overlap between the predicted anomalous region $P$ and the ground truth anomalous region $G$ at a given threshold $t$: $\text{IoU}=|P\cap G|/|P\cup G|=\text{TP}/(\text{TP}+\text{FP}+\text{FN})$.
  • Per-Region Overlap (PRO) [12] Averages the detection rate over each ground truth connected component, giving equal weight to small and large anomalies. $\text{PRO}=\frac{1}{N}\sum_{i}\sum_{j}\frac{|P_{i}\cap G_{i,j}|}{|G_{i,j}|}$, where $N$ is the total number of GT components, $P_{i}$ is the prediction for image $i$, and $G_{i,j}$ is the $j$-th ground truth component in image $i$.
  • Track-Based Detection Rate (TBDR) [169] (Video) Measures the fraction of anomalous object tracks correctly detected (a track is detected if $\alpha$ fraction of its GT regions are detected with IoU $\geq\beta$). Used with False Positive Rate per frame (FPR).
  • Region-Based Detection Rate (RBDR) [169] (Video): Measures the fraction of all anomalous regions across all frames correctly detected (IoU $\geq\beta$). Also evaluated against FPR.
VI-B Metrics for Evaluating Explanations
Evaluating the quality of explanations themselves is more challenging and less standardized. Key aspects include [131, 17]:
  • Faithfulness (Fidelity) Does the explanation accurately reflect the model’s internal reasoning process? For feature attribution methods (like attention or SHAP), one common approach is to measure the change in model output when perturbing or removing features deemed important by the explanation. A faithful explanation should lead to a significant drop in the anomaly score when important features are removed.
  • Plausibility (Sanity) Does the explanation make sense to a human observer? This is often evaluated qualitatively through user inspection. For example, does an attention map highlight the visually anomalous region? Does a textual explanation correctly describe the anomaly? Sanity checks can also involve testing whether explanations change appropriately when model parameters or inputs are randomized [173].
  • Stability (Robustness) Is the explanation robust to small, insignificant perturbations in the input? An unstable explanation that changes drastically with minor input variations may not be reliable.
  • Completeness: Does the explanation capture all the factors influencing the decision?
  • Usefulness/Effectiveness Does the explanation help the user achieve a specific goal, such as debugging the model, understanding its limitations, or making a more informed decision? This often requires task-specific user studies.
  • Consistency with Ground Truth Explanations For datasets like X-MAN [97] that provide ground truth reasons for anomalies (e.g., attribute labels), the consistency between the model’s explanation (e.g., predicted attributes) and the ground truth can be directly measured using metrics like accuracy or F1-score for attribute prediction.
Currently, quantitative evaluation of explainability in visual anomaly detection often relies on proxy tasks. For localization-based explanations (attention maps, error maps), the overlap with ground truth anomaly masks (using IoU or pixel AUROC) is frequently used as a measure of explanation quality, conflating localization performance with explanation faithfulness. For attribute-based explanations, performance on predicting ground truth attributes (if available) serves as a proxy. Developing more direct and standardized metrics for faithfulness, stability, and usefulness remains an open research area, especially for complex visual anomalies. Recent work has also focused on developing practical metrics specifically for explaining abnormal uncertainty in related tasks like object detection [174], which might offer insights for evaluating uncertainty-based explanations in anomaly detection.
VII Future Directions and Open Problems
Based on the reviewed literature and remaining challenges, we outline several promising future directions and open problems in explainable 2D visual anomaly detection.
VII-A Quantifying and Evaluating Explanation Quality
Challenge: While numerous methods provide visual or textual explanations, quantitatively evaluating their quality beyond standard detection/localization metrics (Sec. VI) remains a major hurdle. Current evaluations often rely on qualitative examples or proxy metrics (e.g., localization accuracy for attention maps) that don’t fully capture explanation fidelity or usefulness. Directions:
  • Developing Faithfulness Metrics Adapt and standardize metrics from general XAI research [131] to assess if explanations (e.g., heatmaps from attention or Shapley values [123, 124]) truly reflect the model’s decision process for visual anomalies. This involves measuring performance degradation upon perturbing explained features/regions.
  • Metrics for Textual Explanations For explanations generated by VLMs/LLMs [69, 71], develop metrics beyond semantic similarity to evaluate factual correctness, relevance to the visual anomaly, and absence of hallucination.
  • User Studies Conduct targeted user studies to evaluate the effectiveness and usefulness of different explanation types (e.g., heatmaps vs. attribute lists vs. natural language) for specific tasks (e.g., debugging industrial defects, understanding medical anomalies).[17]
  • Benchmarking Explanation Stability Develop protocols to measure the robustness of explanations to minor input variations, ensuring reliability.
  • Ground Truth Explanations Encourage the creation of more datasets with ground truth explanations (like X-MAN [97]) to enable direct evaluation of attribute-based or reason-based explanations.
VII-B Explaining Diverse Anomaly Detection Paradigms
Challenge: Explainability needs to be integrated into and tailored for various AD settings and techniques. Directions:
  • Self-Supervised Learning (SSL) Current SSL-based AD methods [35, 5, 34, 121, 122] often lack clear explanations for why an anomaly causes the pretext task to fail. Future work should investigate methods to link pretext task performance degradation back to specific anomalous features or temporal inconsistencies.
  • Few-Shot, Zero-Shot, and Open-Set AD Explaining decisions in data-scarce settings (few-shot [175], zero-shot [176]) or when encountering completely novel anomalies (open-set) is crucial. How can models justify flagging something as anomalous when they have limited or no prior exposure? Explanations might involve highlighting deviation from known normality or leveraging external knowledge (e.g., via VLMs).
  • Weakly Supervised AD In settings with only video-level labels [167, 168, 177], developing methods that not only localize anomalies but also provide faithful explanations for the localization remains challenging.
  • Feature-Based Methods For methods operating on pre-trained features [117, 118, 119, 3, 120], develop techniques to translate feature-level anomaly explanations (e.g., deviations in feature space) into interpretable concepts in the input space.
VII-C Context-Aware and Reasoning-Based Explanations
Challenge: Many anomalies are context-dependent (e.g., a person running is normal on a track but abnormal in a library). Current explanations often focus on local features and struggle with contextual reasoning. Directions:
  • Scene-Context Modeling Develop explainable models that explicitly incorporate scene context (e.g., using scene graphs [89], knowledge graphs [115], or attention mechanisms that consider background) to justify why an object or event is anomalous in that specific context. The NWPU Campus dataset [172] is a good testbed for this.
  • Causal Explanations Move beyond correlational explanations (e.g., attention maps) towards causal explanations [88] that identify the underlying causes of anomalies, especially for complex events in videos.
  • Integrating Commonsense Knowledge Leverage external commonsense knowledge bases or the implicit knowledge in foundation models to provide more human-like reasoning for why certain situations are anomalous.
VII-D Explainable Foundation Models for Visual Anomaly Detection
Challenge: While foundation models (VLMs/LLMs) [178, 179, 180, 181] show promise for zero-shot AD and generating textual explanations [69, 71, 72], ensuring the faithfulness and controllability of these explanations is crucial. Applying them effectively to VAD also requires addressing temporal modeling. Directions:
  • Faithful Textual Explanations Develop methods to ensure that the natural language explanations generated by LLMs accurately reflect the visual evidence and the model’s internal anomaly detection process, mitigating hallucination.
  • Interactive Explanations Explore interactive explanation systems where users can query the model (e.g., ”Why is this region anomalous?”, ”What would need to change for this to be normal?”) to gain deeper understanding.
  • Temporal Reasoning with Foundation Models Adapt foundation models for robust temporal reasoning in VAD, enabling explanations that capture complex event dynamics, object interactions, and motion anomalies [132, 90, 133].
  • Efficiency and Controllability Improve the efficiency of using large foundation models for AD and develop ways to control the level of detail and perspective of the generated explanations.
VII-E Real-World Considerations: Efficiency and Robustness
Challenge: Practical deployment requires explainable AD systems to be efficient and robust. Directions:
  • Efficient Explanations Develop explanation techniques (especially for complex methods like Shapley values or reasoning-based approaches) that are computationally efficient enough for real-time applications.
  • Robustness of Explanations Evaluate and improve the robustness of explanations against adversarial attacks specifically designed to mislead the explanation module while potentially leaving the anomaly detection itself unaffected.
  • Domain Adaptation and Generalization Ensure that explainable AD methods generalize well to new domains and that the explanations remain meaningful and consistent across different operational conditions.